I need to solve this with a equation.
The sport shop sells bicycles and tricycles and often has a selction of them sitting out for customers to try. While I was looking around, my son counted 70 wheels and 29 handlebars. How many bicycles and how many tricyles were there?

Answer by scott8148(6628)

70=2b+3t ... b+t=29 ... b=29-t ... 70=2(29-t)+3t ... 70=58-2t+3t ... 12=t ... b+12=29 ... b=17